OpenStudy (anonymous):

How would i solve this : 2x-4y=12

OpenStudy (anonymous):

Does the problem specify what they are looking for? Are there additional details for the problem? This equation cannot be solved, it can be rearranged however, to match a certain form.

OpenStudy (anonymous):

the question said " Draw a complete graph of the line 2x-4y=12 using any two points."

OpenStudy (anonymous):

Ahhhh :) Well, I can help you with that as well!

OpenStudy (anonymous):

First I need to know, are you familiar with the form y = mx + b?

OpenStudy (anonymous):

yes, i am i believe that is the formula for slope intercept

OpenStudy (anonymous):

Correct, it is the formula for a linear equation.

OpenStudy (anonymous):

Can I get you to put 2x-4y=12 into y=mx+b form?

OpenStudy (anonymous):

Do you know how to proceed?

OpenStudy (anonymous):

would it be : 4y=2x-12

OpenStudy (anonymous):

Excellent! That is absolutely correct. Now to make this easier on us, can you get rid of the coefficient (the constant in front) of y?

OpenStudy (anonymous):

That is to say, can you get rid of that 4 in front of the y?

OpenStudy (anonymous):